OpenStudy (anonymous):

Evaluate the function at the given values for x: f(x) = log2 x + 1 Write your answer in ordered pair form in the blank. Ex: (1/4, ), (1, ), (4, )

OpenStudy (anonymous):

\[f(x)=\log_2(x)+1\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f(1)=\log_2(1)+1=0+1=2\]

OpenStudy (anonymous):

the reason \[\log_2(1)=0\] is because \[2^0=1\]

OpenStudy (anonymous):

\[\log_2(4)+1=2+1=3\]

OpenStudy (anonymous):

and again \[\log_2(4)=2\] because \[2^2=4\]

OpenStudy (anonymous):

finally \[\log_2(\frac{1}{4})+1=-2+1=-1\]

OpenStudy (anonymous):

because \[2^{-2}=\frac{1}{2^2}=\frac{1}{4}\]

OpenStudy (anonymous):

I'm sorry, I dont understand where the ordered pair is.

OpenStudy (anonymous):

is it (1/4, -1)?